OpenStudy (zeesbrat3):

If f(x) = |(x2 − 4)(x2 + 2)|, how many numbers in the interval [0, 1] satisfy the conclusion of the Mean Value Theorem?

OpenStudy (zeesbrat3):

@Astrophysics @ganeshie8 @Michele_Laino

OpenStudy (zeesbrat3):

help..

OpenStudy (michele_laino):

I'm pondering...

OpenStudy (zeesbrat3):

Oh, sorry

OpenStudy (michele_laino):

we have to note that we have the subsequent inequality: \[\Large {x^2} - 4 < 0,\quad {\text{if }} - 2 < x < 2\]

OpenStudy (michele_laino):

am I right?

OpenStudy (zeesbrat3):

honestly, im confused

OpenStudy (michele_laino):

why?

OpenStudy (michele_laino):

it is a simple inequality

OpenStudy (zeesbrat3):

no i understand the inequality, im confused by how you got it

OpenStudy (michele_laino):

ok! I wrote that inequality, because inside the interval [0,1] your function can be rewritten as follows: \[\Large f\left( x \right) = - \left( {{x^2} - 4} \right)\left( {{x^2} + 2} \right) = \left( {4 - {x^2}} \right)\left( {{x^2} + 2} \right)\]

OpenStudy (zeesbrat3):

otherwise any other would be out of bounds

OpenStudy (michele_laino):

no, it is a step in order to get rid the absolute value symbol

OpenStudy (zeesbrat3):

oh, okay.

OpenStudy (michele_laino):

now, we can apply the theorem af mean value

OpenStudy (michele_laino):

so, we have to compute this: \[\Large \frac{{f\left( 1 \right) - f\left( 0 \right)}}{{1 - 0}} = \frac{{3 \times 3 - 4 \times 2}}{1} = ...\]

OpenStudy (zeesbrat3):

\[f'(c) = \frac{ f(b) - f(a) }{ b - a }\]

OpenStudy (zeesbrat3):

1

OpenStudy (michele_laino):

are you sure?

OpenStudy (michele_laino):

\[\large \frac{{f\left( 1 \right) - f\left( 0 \right)}}{{1 - 0}} = \frac{{3 \times 3 - 4 \times 2}}{1} = \frac{{6 - 8}}{1} = ...?\]

OpenStudy (zeesbrat3):

isnt 3x3 9?

OpenStudy (michele_laino):

sorry I have made a big error, you are right

OpenStudy (zeesbrat3):

it's okay :)

OpenStudy (michele_laino):

sorry again!!

OpenStudy (zeesbrat3):

no apologies necessary! thank you for your help

OpenStudy (michele_laino):

next we have to compute the first derivative at a generic point x, namely we have to compute the first derivative of this function: \[\Large f\left( x \right) = \left( {4 - {x^2}} \right)\left( {{x^2} + 2} \right)\]

OpenStudy (michele_laino):

what do you get?

OpenStudy (zeesbrat3):

\[-4x(x^2-1)\]

OpenStudy (michele_laino):

correct!

OpenStudy (michele_laino):

then we have to solve this equation: \[ \Large - 4x\left( {{x^2} - 1} \right) = 1\]

OpenStudy (michele_laino):

it is not a simple equation

OpenStudy (zeesbrat3):

\[x^3 = -\frac{ 1 }{ 2 }\] ?

OpenStudy (michele_laino):

how did you do to write that expression?

OpenStudy (zeesbrat3):

\[-4x(x^2-1) = 1\]\[\frac{ 1 }{ -4x } = x^2\] \[\frac{ 1 }{ -4x } +1 = x^2\] \[1+1 = -4x^3\] \[- \frac{ 1 }{ 2 } = x^3\]

OpenStudy (michele_laino):

I don't understand this step: \[\frac{1}{{ - 4x}} = {x^2}\]

OpenStudy (michele_laino):

anyway it is not a solution of our equation

OpenStudy (zeesbrat3):

Sorry, i forgot the -1

OpenStudy (zeesbrat3):

How would you get the proper solution?

OpenStudy (michele_laino):

I questioned wolfram alpha, here are the solutions:

OpenStudy (zeesbrat3):

So, 3 solutions in the interval

OpenStudy (michele_laino):

we have two solutions, inside the interval [0,1]

OpenStudy (zeesbrat3):

Whoops, my bad! I looked outside the interval

OpenStudy (michele_laino):

do you see them?

OpenStudy (zeesbrat3):

Yes. Since one of them is in the negatives, it is not in the interval and does not matter to answer the question

OpenStudy (michele_laino):

yes! that's right!

OpenStudy (zeesbrat3):

Thank you!

OpenStudy (michele_laino):

:)

OpenStudy (zeesbrat3):

I tried doing the same thing on this problem, but I keep getting no solutions: A particle moves along the x-axis with position function s(t) = xe^x. How many times in the interval [−5, 5] is the velocity equal to 0?

OpenStudy (loser66):

velocity = (s(t))' take derivative and let it =0 to solve for x .